# Differential Privacy as a Perk: Federated Learning over Multiple-Access Fading Channels with a Multi-Antenna Base Station

### v2

@@ V Experiments
In this section, we present numerical experiments to corroborate our theoretical analyses for privacy and convergence of the AirFL-DP training protocol, as well as verifying in practical settings the range of prescribed privacy level and received SNR where DP can be gained as a perk.
Figure 3: The evolution of the (a) testing accuracy and (b) training loss versus the global communication rounds $T$ with different privacy budgets $\tilde{\epsilon}$.
Figure 4: The evolution of the (a) testing accuracy and (b) training loss versus the privacy budget $\tilde{\epsilon}$. Note that benchmarks relate to the case of $\tilde{\epsilon}\to\infty$ and are drawn in the same figure for illustration purposes only.
@@ V-A Experimental Setting
We evaluate our protocol on the Fashion-MNIST dataset [41] for a $10$-class image classification task. The dataset, comprising $60,000$ training and $10,000$ test images (each $28\times 28$ pixels), is partitioned in an i.i.d. manner across the clients. Each client holds a disjoint subset of the training data, equally balanced across all classes. The clients train a common CNN model which consists of three convolutional layers and a softmax output layer, totaling $d=582,026$ parameters. All reported results are averaged over $10$ Monte Carlo trials to ensure statistical significance.
Unless otherwise specified, we use the following parameters for our experiments. We simulate the FL system with $n=50$ WDs and a BS equipped with $m=100$ antennas. In each of the $T=50$ communication rounds, all clients participate (client sampling rate $r=1$) and perform $Q=5$ local SGD steps with a mini-batch size of $10$. The learning rate is set to $\eta=0.005$, the clipping threshold $c=\sqrt{0.012d}$, and the transmit power is $P=2$ mW for all devices. For the DP concern, we set the parameters to $\tilde{\epsilon}\triangleq\epsilon\mathord{\left/\right.\kern-1.2pt}\sqrt{d}=0.1$ and $\delta=10^{-5}$.
We adopt a block-flat Rayleigh fading channel, i.e., $\boldsymbol{h}_{i}^{(t)}\sim\mathcal{CN}(\boldsymbol{0},\Lambda\boldsymbol{I})$. The path loss coefficient $\Lambda$ is determined by a simplified path loss model, i.e., $\Lambda=(\frac{c_{l}}{4\pi f_{c}r_{i}})^{2}$, where $c_{l}$ denotes the speed of light, $f_{c}=2.4$ GHz is the central frequency, and the device-to-BS distance $r_{i}\sim 1000\times\sqrt{\mathcal{U}(0,1)}$. Furthermore, the power spectrum density is set to $N_{0}=-173$ dBm/Hz over a bandwidth of $20$ MHz, resulting in the channel noise scale $\sigma^{2}\approx-100$ dBm.
We consider the following benchmarks:
@@
    AirFL-ZF: This scheme employs ZF beamforming (36) to maximize performance over the fading channel without considering the DP constraint.
@@ V-B Experimental Analysis
Figure 5: The evolution of the (a) testing accuracy and (b) training loss versus the worst-case SNR, $P\Lambda_{\min}/\sigma^{2}$, where $\Lambda_{\min}=(c_{l}\mathord{\left/\right.\kern-1.2pt}(4\pi f_{c}r_{\max}))^{2}$ and $r_{\max}=1000$. The SNR is varied by adjusting the transmit power $P$.
In Fig. 3, we report the test accuracy and training loss versus the communication round $T$ with different privacy budgets $\tilde{\epsilon}$, respectively. The results reveal the inherent privacy-convergence trade-off, where stricter privacy guarantees result in lower test accuracy and higher training loss. Notably, as the privacy budget is relaxed (e.g., to $\tilde{\epsilon}=0.13$, the performance gap between the AirFL-DP and AirFL-ZF significantly decreases, indicating that AirFL-DP achieves privacy protection for free without performance degradation for a relatively strict privacy level. Furthermore, we observe that the performance of AirFL-ZF is highly comparable to that of FL w/ clipping operating on a perfect channel. It confirms that the ZF solution can effectively mitigate the wireless channel impairments, demonstrating its optimality.
Fig. 4 shows the test accuracy and training loss versus the privacy budget $\tilde{\epsilon}$, where a small $\tilde{\epsilon}$ corresponds to a more stringent privacy requirement. The performance of AirFL-DP sharply approaches the benchmarks as $\tilde{\epsilon}$ increases, demonstrating that AirFL-DP provides free DP guarantees without sacrificing the performance in a wide range of privacy budgets, thanks to the tighter privacy bound and optimal transceiver design.
Finally, Fig. 5 validates our theoretical analysis on the impact of the SNR. This result reveals a distinct threshold effect, as shown in (52). In the low-SNR regime, the performance of AirFL-DP is identical to that of AirFL-ZF, while it performs slightly poorer than AirFL-ZF as SNR increases, which coincides with the observation in Sec. IV.
@@ VI Conclusions and Discussions
Lastly, we briefly discussed the solution to the general problem $(\text{P0}^{\prime})$ with a finite $D$, where the privacy constraint includes the term $\min\{\sum_{t}\phi_{t},\Phi\}$. As mentioned, this problem remains tractable and can be solved using the same principles as the simplified case. A standard approach to handle the $\min$ operator is to introduce an auxiliary constraint, which can be incorporated into the Lagrangian function via an additional dual variable. This, in turn, necessitates a subsequent case analysis based on whether the new multiplier is active or inactive. While the resulting derivation becomes more algebraically involved, the core solution approach—rooted in the KKT conditions—remains unchanged. For brevity, we omitted the full derivation, as it follows the same logical steps and offers no new conceptual insights.
Appendix A Important Lemmas
Lemma A.1 (Equivalent Definitions of $\infty$-Wasserstein Distance [30])
The following are equivalent for any distributions $\mathbb{P}_{\boldsymbol{\mu}}$ and $\mathbb{P}_{\boldsymbol{\nu}}$:
  1. 1.
    $W_{\infty}(\mathbb{P}_{\boldsymbol{\mu}},\mathbb{P}_{\boldsymbol{\nu}})\leq s$.
  2. 2.
    There exist jointly distributed random variables $(\boldsymbol{u},\boldsymbol{v})$ such that $\boldsymbol{u}\sim\mathbb{P}_{\boldsymbol{\mu}}$, $\boldsymbol{v}\sim\mathbb{P}_{\boldsymbol{\nu}}$, and $\operatorname{Pr}[\|\boldsymbol{u}-\boldsymbol{v}\|\leq s]=1$.
  3. 3.
    There exist jointly distributed random variables $(\boldsymbol{u},\boldsymbol{w})$ such that $\boldsymbol{u}\sim\mathbb{P}_{\boldsymbol{\mu}}$, $\boldsymbol{u}+\boldsymbol{w}\sim\mathbb{P}_{\boldsymbol{\nu}}$, and $\operatorname{Pr}[\|\boldsymbol{w}\|\leq s]=1$.
Lemma A.2 (Non-Expansiveness of Clipping)
For any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$, we have
$\|\operatorname{clip}_{c}(\boldsymbol{x})-\operatorname{clip}_{c}(\boldsymbol{y})\|\leq\|\boldsymbol{x}-\boldsymbol{y}\|.$ (54)
Lemma A.3 (Partial Convexity in the Second Argument [42])
Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}_{1}$, and $\boldsymbol{\nu}_{2}$ be three random variables. Then, for any $\alpha\geq 0$, the Rényi divergence is convex in its second argument, that is, the following inequality holds
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||(1-\lambda)\mathbb{P}_{\boldsymbol{\nu}_{1}}+\lambda\mathbb{P}_{\boldsymbol{\nu}_{2}})$ (55)
$\leq(1-\lambda)\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}_{1}})+\lambda\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}_{2}}),$
for any $0\leq\lambda\leq 1$.
Lemma A.4 (Post-Processing Inequality [42])
Let $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ be two random variables. Then, for any (possibly random) function $\psi$ and $\alpha\geq 0$, we have
$\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi(\boldsymbol{\nu})})\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}}).$ (56)
Lemma A.5 (Strong Composition for Rényi Divergence [33])
Let $\mathcal{M}=\left[\mathcal{M}_{1}(\mathcal{D}),\mathcal{M}_{2}\left(\mathcal{D}\right)\right]$ be a sequence of two randomized mechanisms. Then, for any $\alpha>1$, we have
$\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}(\mathcal{D})}||\mathbb{P}_{\mathcal{M}(\mathcal{D}^{\prime})})\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}_{1}(\mathcal{D})}||\mathbb{P}_{\mathcal{M}_{1}(\mathcal{D}^{\prime})})$ (57)
$+\sup_{\boldsymbol{v}}\mathcal{D}_{\alpha}(\mathbb{P}_{\mathcal{M}_{2}(\mathcal{D})|\mathcal{M}_{1}(\mathcal{D})=\boldsymbol{v}}||\mathbb{P}_{\mathcal{M}_{2}(\mathcal{D}^{\prime})|\mathcal{M}_{1}(\mathcal{D}^{\prime})=\boldsymbol{v}}),$
where $\mathcal{D}$ and $\mathcal{D}^{\prime}$ are two adjacent datasets.
Lemma A.6
For any vector $\boldsymbol{x}$ and $\boldsymbol{y}$, it holds that
$\boldsymbol{x}^{T}\boldsymbol{y}=-\frac{1}{2}\|\boldsymbol{x}\|^{2}-\frac{1}{2}\|\boldsymbol{y}\|^{2}+\frac{1}{2}\|\boldsymbol{x}+\boldsymbol{y}\|^{2}.$ (58)
Lemma A.7
For any $a,b\in\mathbb{R}$, it holds that
$(a+b)^{2}\leq 2a^{2}+2b^{2}.$ (59)
Lemma A.8
For any learning rate $\eta$ satisfying $\eta\leq\frac{1}{8QL}$, under Assumptions 1 and 4-5, we can bound the drift for any $q\in\{0,\ldots,Q-1\}$ as
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left\|\boldsymbol{\theta}_{i}^{(t,q)}-\boldsymbol{\theta}^{(t)}\right\|^{2}$ (60)
[…]
Appendix B Proof of Lemma III.3
@@ Appendix A Important Lemmas
where the first step is by the triangle inequality, Assumption 1, and Lemma A.2, and the second step is by the definition of $\boldsymbol{\mu}^{\prime}$. When round $t$ involves $Q$ successive local SGD updates, the overall mapping $\psi^{(t)}$ can be viewed as the composition of $Q$ such one-step mappings. Then, the $Q$-step bound immediately follows as
@@
$\leq(1-r)\mathcal{D}_{\alpha}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu}^{\prime})}||\mathbb{P}_{\psi^{(t)}(\boldsymbol{\nu})})+r\mathcal{D}_{\alpha}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu}^{\prime})}||\mathbb{P}_{\psi^{(t)\prime\prime}(\boldsymbol{\nu})})$
@@
where the first step is by Definition III.2, the second step is due to device sampling with $\psi^{\prime\prime}(\cdot)\stackrel{{\scriptstyle\Delta}}{{=}}\psi^{\prime}(\cdot|i^{*}\in\mathcal{I}^{(t)})$, the third step is by Lemma A.3, and the last step is by Lemma A.4.
@@
where the first step is by Lemma A.4, the second step is by the Lemma A.5, and the last step is by the well-known result $\mathcal{D}_{\alpha}(\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{d})\|\mathcal{N}(\boldsymbol{u},\sigma^{2}\boldsymbol{I}_{d}))=\alpha\|\boldsymbol{u}\|_{2}^{2}/2\sigma^{2}$. Hence,
@@
Appendix C Proof of Proposition III.1
@@
$=\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\theta}^{(t)})+\boldsymbol{\varsigma}^{(t)}}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\theta}^{\prime(t)})+\boldsymbol{\varsigma}^{(t)}})$
@@
$=\mathcal{D}_{\alpha}^{((1+\kappa_{\max})^{Q}z_{t})}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\theta}^{(t)})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\theta}^{\prime(t)})})+\frac{\alpha a_{t}^{2}r^{2}n^{2}}{2(1-\beta^{(t)})W^{(t)}}$
@@
where the second step is by Lemma III.2, the fourth step is by the Definition III.2, and the last step is by Lemma III.3. By repeating the induction from $T$ to $\tau$, we can obtain
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}^{(T)}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime(T)}})\!\leq\!\sum_{t=\tau}^{T-1}\frac{2\alpha rc^{2}\phi_{t}}{\beta^{(t)}\sigma^{2}}\!+\!\frac{\alpha(1+\kappa_{\max})^{2Q}\Xi_{\tau}^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(\tau)})\|\boldsymbol{w}^{(\tau)}\|^{2}}$ (71)
@@
Appendix D Proof of Corollary III.1

### v3

@@ V Experiments
In this section, we present numerical experiments to corroborate our theoretical analyses for privacy and convergence of the AirFL-DP training protocol, as well as verification in a varied range of prescribed privacy levels and received SNR.
Figure 5: The evolution of the testing accuracy versus the privacy budget $\tilde{\epsilon}$. Note that benchmarks relate to the case of $\tilde{\epsilon}\to\infty$ and are drawn in the same figure for illustration purposes only.
@@ V-A Experimental Setting
We evaluate our protocol on the Fashion-MNIST dataset [35] for a $10$-class image classification task. The dataset, comprising $60,000$ training and $10,000$ test images (each $28\times 28$ pixels), is partitioned in an i.i.d. manner across the clients. Each client holds a disjoint subset of the training data, equally balanced across all classes. The clients train a common CNN model which consists of three convolutional layers and a softmax output layer, totaling $d=582,026$ parameters. All reported results are averaged over 1010-3030 Monte Carlo trials to ensure statistical significance.
Unless otherwise specified, we use the following parameters for our experiments. We simulate an FL system with $n=50$ WDs and a BS equipped with $m=100$ antennas. In each of the $T=50$ communication rounds, clients with sampling rate $r=0.9$ perform $Q=5$ local SGD steps with a mini-batch size of $10$. Other parameters are set as follows: the learning rate $\eta=0.005$, the clipping threshold $c=\sqrt{0.012d}$, the transmit power $P=2$ mW for all devices, and the DP parameters $\tilde{\epsilon}\triangleq\epsilon\mathord{\left/\right.\kern-1.2pt}\sqrt{d}=0.1$ and $\delta=10^{-5}$.
We adopt a block-flat Rayleigh fading channel, i.e., $\boldsymbol{h}_{i}^{(t)}\sim\mathcal{CN}(\boldsymbol{0},\Lambda\boldsymbol{I})$. The path loss coefficient is determined by a simplified path loss model, i.e., $\Lambda=(\frac{c_{l}}{4\pi f_{c}r_{i}})^{2}$, where $c_{l}$ denotes the speed of light, $f_{c}=2.4$ GHz is the central frequency, and the device-to-BS distance $r_{i}\sim 1000\times\sqrt{\mathcal{U}(0,1)}$ m. Furthermore, the power spectrum density is set to $N_{0}=-173$ dBm/Hz over a bandwidth of $20$ MHz, resulting in the channel noise scale $\sigma^{2}\approx-100$ dBm.
We compared the performance of AirFL-DP to the following benchmarks:
@@
    AirFL-MIMO: This scheme employs the beamforming scheme [40] to maximize convergence performance over the fading channel without considering the DP constraint.
@@ V-B Experimental Analysis
Figure 6: The evolution of the testing accuracy versus the worst-case SNR, $P\Lambda_{\min}/\sigma^{2}$, where $\Lambda_{\min}=(c_{l}\mathord{\left/\right.\kern-1.2pt}(4\pi f_{c}r_{\max}))^{2}$ and $r_{\max}=1000$. The SNR is varied by adjusting the transmit power $P$.
In Fig. 3, we report the test accuracy and training loss versus the communication round $T$ with different privacy budgets $\tilde{\epsilon}$, respectively. The results reveal the inherent privacy-convergence trade-off, where stricter privacy guarantees result in lower test accuracy and higher training loss. Notably, as the privacy budget is relaxed, e.g., to $\tilde{\epsilon}=0.13$, the performance gap between the AirFL-DP and AirFL-MIMO significantly decreases, indicating that AirFL-DP achieves privacy protection for free for a relatively strict privacy level. Furthermore, we observe that the performance of AirFL-MIMO is highly comparable to that of FL w/ clipping operating on a perfect channel. It confirms that this solution can effectively mitigate the wireless channel impairments, demonstrating its optimality.
We then evaluate the proposed method under the non-i.i.d. scenario in Fig. 4. We adopt the commonly used non-i.i.d. setting, in which different client observes data from different subsets of label classes. Following standard practice [28], the dataset is first grouped by labels, and each client is assigned data from only $k$ classes. A smaller $k$ leads to stronger heterogeneity. Here, we set $k=2$, resulting in highly skewed label distributions across clients. Compared to Fig. 3(b) under the i.i.d. settings, due to strong label distribution skew, the overall accuracy in Fig. 4(b) is slightly lower, which aligns with expectations.
Fig. 5 illustrates the test accuracy versus the privacy budget $\tilde{\epsilon}$, where a small $\tilde{\epsilon}$ corresponds to a more stringent privacy requirement. The performance of AirFL-DP drastically climbs up to the benchmarks as $\tilde{\epsilon}$ increases, demonstrating that AirFL-DP provides free DP guarantees without much compromise to performance in a wide range of privacy budgets, thanks to the tighter privacy bound and optimal transceiver design.
Finally, Fig. 6 validates the impact of the SNR on the performance, which we observe from the theoretical analysis. This result reveals a distinct thresholding effect, as shown in (51). In the low-SNR regime, the performance of AirFL-DP is identical to that of AirFL-MIMO, while it performs slightly poorer than AirFL-MIMO as SNR increases, which aligns with the conclusion drawn in Remark IV.2.
@@ VI Conclusions and Discussions
Appendix A Proof of Lemma III.3
@@
where the first step is by the triangle inequality, Assumption 1, and the non-expansiveness of clipping [3], and the second step is by the definition of $\boldsymbol{\mu}^{\prime}$. When round $t$ involves $Q$ successive local SGD updates, the overall mapping $\psi^{(t)}$ can be viewed as the composition of $Q$ such one-step mappings. Then, the $Q$-step bound immediately follows as
@@
where the first step is by Definition III.2, the second step is due to device sampling with $\psi^{\prime\prime}(\cdot)\stackrel{{\scriptstyle\Delta}}{{=}}\psi^{\prime}(\cdot|i^{*}\in\mathcal{I}^{(t)})$, the last step is by the partial convexity [30, Theorem 12] and post-processing inequality [30, Theorem 9].
@@
where the first step is by the post-processing inequality, the second step is by the strong composition of RDP [23, Proposition 1], and the last step is by the well-known result $\mathcal{D}_{\alpha}(\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{d})\|\mathcal{N}(\boldsymbol{u},\sigma^{2}\boldsymbol{I}_{d}))=\alpha\|\boldsymbol{u}\|_{2}^{2}/2\sigma^{2}$. Hence,
@@
Appendix B Proof of Proposition III.1
@@
where the first step is by Lemma III.2, the second step is by the Definition III.2, and the last step is by Lemma III.3. By repeating the induction from $T$ to $\tau$, we can obtain
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}^{(T)}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime(T)}})\!\leq\!\sum_{t=\tau}^{T-1}\frac{2\alpha rc^{2}\phi_{t}}{\beta^{(t)}\sigma^{2}}\!+\!\frac{\alpha(1+\kappa_{\max})^{2Q}\Xi_{\tau}^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(\tau)})\|\boldsymbol{w}^{(\tau)}\|^{2}},$ (63)
@@
Appendix C Proof of Corollary III.1